OpenStudy (anonymous):

Not a clue. The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (jdoe0001):

in short \(\huge d = ut + 0.5 at^2\) solve for "a"

OpenStudy (anonymous):

Yes, but I don't understand how to do it at all.

OpenStudy (anonymous):

to solve an equation for some variable, you are always allowed to apply the same operation on both sides. For example, you can subtract ut from both sides: d-ut = 0.5 at^2

OpenStudy (anonymous):

Would I then also subtract t as im solving for a? I've not understood the lesson at all so im winging it :d.

OpenStudy (anonymous):

you can't get the t away by subtraction, because it was not "tied" to a by addition. In the first part, the ut WAS tied by addition, so we used subtraction to cancel it. This time it's MULTIPLIED with a, so we have to DIVIDE to split it from a.

OpenStudy (anonymous):

First we had\[d=ut+0.5 a t^{2}\] which is nothing else than \[d=ut+0.5 \times a \times t^{2}\]. Then, subtract ut from both sides: \[d-ut=0.5 \times a \times t^{2}\] the last step is to divide by 0.5 t^2 on BOTH SIDES: \[\frac{ d-ut }{ 0.5 t^{2} } = a\]